OpenStudy (anonymous):

A train leaves Lexington for Indianapolis, 200 miles awa y, at 1:00 PM and averages 60 miles per hour. A second train travelling on an adjacent track leaves I ndianapolis for Lexington at 3:30 PM and averages 40 miles per hour. At what time will the trains me et? (Round to the nearest minute.)

OpenStudy (raden):

60t + 40(t - 2.5) = 200 solve for t

OpenStudy (raden):

did u get it ?

OpenStudy (anonymous):

I found 3h as answer.

OpenStudy (anonymous):

I thought it would be something like (60t)^2 + (( 40 (t-2.5))^2= 200^2 ?